# WindGP: Efficient Graph Partitioning on Heterogenous Machines

### 01 Mar 2024 v1

@@ 5.2. Evaluation by the metric $TC$
Analysis of optimization techniques. Recall that WindGP propose three novel techniques: preprocessing capacity, best-first search, post-processing. To evaluate the efficiency of each technique in WindGP, we conduct the experiments on six real graphs and depict the results in Figure 8. Let WindGP be the entire solution, ${}^{+}$ denotes the solution that removes post-processing from WindGP. Similarly, $*{}^{*}$ denotes the solution that removes best-first search from ${}^{+}$, while ${}^{-}$ is the naive solution without all optimization techniques. In fact, ${}^{-}$ iteratively explores the graph to form a partition as long as the machine memory permits. Its expansion scheme is similar to NE (Zhang et al., 2017), which tends to form a subgraph with high cohesion by connectivity-based expansion. In order to view the details of all partitions precisely, the histograms of the costs of different partitions on CP and LJ are also listed in Figure 9 and Figure 10, respectively.
@@
Obviously, $TC$ of $*{}^{*}$ is much lower than ${}^{-}$, and the speedup is 5×5$\times$5 × on TW, 4.5×4.5$\times$4.5 × on CO, $\times3$ on LJ, 3.8×3.8$\times$3.8 × on PO, 7.5×7.5$\times$7.5 × on CP, and $\times8$ on RN, respectively. This prominent improvement comes from the pre-computed capacity of edges in each machine. Different from the naive solution that only considers the upper bound of machine memory, our strategy combines the computation cost and the memory size of all machines to generate a nearly best plan that can balance the computation cost across these heterogeneous machines (as shown in Figure 9(b) and Figure 10(b)). Note that traditional partition methods use $\frac{|E|}{p}$ as the capacity of edges, which can not be used in heterogeneous scenarios because the memory of some machines may be not enough while other machines may have memory size larger than $\frac{|E|}{p}$. Besides, the traditional threshold does not distinguish the difference of node/edge computation cost between heterogeneous machines. Recall that in Section 3.2 we deduce the error bound $\frac{p^{2}}{|E|}$, and the experiment on small graphs with hundreds of edges shows that the difference between our solution and the optimal is always within 5%, which can be further refined by the post-processing. According to the theoretical bound, on larger graphs the error should be much smaller.
@@
The second technique (i.e., the best-first search) further boosts the performance by optimizing the expansion process, leading to >1.1×>1.1$\times$> 1.1 × on TW, CO and LJ. The performance gain is mainly acquired by lowering the total communication cost (as shown in Figure 9(c) and Figure 10(c)). The degree balanced generation as well as border generation can help reduce the number of cross-machine vertices and improve the subgraph cohesion. However, on low-degree graphs (CP and RN), the speedup is rather limited (only 1.04$\sim$ 1.04$\times$ 1.04 ×) due to the small percentage of communication cost. From Figure 9, the percentage of communication cost is $<30\%$ on sparse graphs (i.e., graphs with low average degree $\frac{|E|}{|V|}$), while it is $>50\%$ on LJ and CO. Especially, the effect is tiny on RN because it is mesh-like graph and its structure is naturally balanced.
@@
Finally, the post-processing technique (subgraph-local search) brings >1.15×>1.15$\times$> 1.15 × speedup by flattening the total cost between all machines. Note that the previous two techniques balance the computation cost and the communication cost respectively, however, they do not balance the total cost well. In fact, the effect of best-first search is strictly restricted because it can not disturb the balance of computation cost. Besides, during expansion the information is not complete (e.g., the first partition has no communication cost), which further drags down the performance. The third technique breaks this restriction by allowing the computation cost to be imbalanced between heterogeneous machines, as long as the total cost drops. Based on the global view with complete information, it moves or swaps edges between machines to lower the highest total cost.
@@
Comparison with counterparts. Figure 12 lists the comparison of WindGP (the entire solution) with the state-of-the-art algorithms on heterogeneous machines. As reported in previous work (Zhang et al., 2017), METIS can not partition too large graphs like TW due to its prohibitive memory occupation. Among four counterparts, NE performs the best on most datasets except for CP and RN. The two low-degree graphs have larger computation cost and lower communication cost, thus the expansion scheme of NE can not optimize a lot. In contrast, METIS and HDRF perform better on CP and RN due to their elaborated consideration of computation cost in the partition objective. As for EBV solution, though it is carefully designed to reduce the effect of skew on power-law graphs, its objective is not equivalent to the maximum total cost and it can not be well adapted to heterogeneous cases.
@@
Due to the lightweight preprocessing and limited iterations in post-precessing, the complexity of WindGP is linear to the graph size and machine number, which is similar to NE. The experiments also show that their running time is in the same scale, while WindGP is 11% slower than NE, which is acceptable in real cases as partitioning is not time sensitive. Besides, the performance of WindGP on homogeneous clusters of virtual machines is not worser than others, as Section 2.1 proves that the $TC$ metric is equivalent to load balance and $RF$ in homogeneous scenarios.
@@ 5.3. Scalability Test
Scalability with the number of machine types. In previous experiments, we use two kinds of machines by default. Here we vary the number of machine types from one to six and test the performance on LJ with 30 machines. Note that the 2-type setting here is different from the default setting. Compared to the single-type setting (i.e., homogeneous cases where all machines are normal machine), 2-type setting transforms 5 machines to slightly bigger machine with larger computation/communication cost as well as larger memory size.
Generally, $TC$ increases as the number of machine types grows, i.e., the homogeneous cases achieve the minimum $TC$ for all solutions. Note that NE performs extremely good on homogeneous cases, even better than WindGP. However, both NE and EBV fail to adapt to heterogeneous cases and their $TC$ rises sharply when the number grows. Comparatively, WindGP achieves the slowest growth due to its flexible preprocessing technique, which can pre-computes appropriate edge capacities for all machine settings.
@@ 5.4. Evaluation of Distributed Graph Algorithms
To evaluate the real performance of distributed graph computing, we choose PageRank (Pretto, 2002) and Single-Source Shortest Path ((Forster and Nanongkai, 2017)) and perform the test on Plato system (Tencent, 2019). Modern machines have large memory and computing power, thus graphs with billions of edges are used in this section. Except for TW, referring to (Zhong et al., 2023) and (Zeng et al., 2022c), we add DB (0.23B nodes and 1.1B edges), FR (65M nodes, 1.8B edges), and YH (0.41B nodes and 2.8B edges). Similar to (Zhang et al., 2017), we use nine machines connected with Gigabit Ethernet, and $\frac{1}{3}$ of them are super machines. The computation cost and communication cost of each machine ($C_{i}^{node}$, $C_{i}^{edge}$ and $C_{i}^{com}$) is estimated according to the number of cores as well as the bandwidth of network. Table 5 and 6 list the comparison of $TC$ and elapsed time (seconds), respectively. Due to the space limit, we only report the performance of HDRF, NE and WindGP.
@@
DB xx xx xx
FR xx xx xx
YH xx xx xx
Obviously, WindGP has the lowest $TC$ value on all datasets, and the effect is rather prominent on enormous skewed graphs (TW and CO). For most cases, a reduction of $TC$ (i.e., the estimated total cost) means a reduction in the running time of distributed graph computing as the partition quality is improved. Though the optimization effect of our techniques is tiny on mesh-like graphs like RN, the running time still drops a little thanks to our sophisticated post-processing technique which considers all kinds of corner cases. The running time of both PageRank and triangle counting is much larger than that on homogeneous machines, because the distributed computing on heterogeneous machines is much more complicated. Super machines have large memory but high computation/communication cost, while the case of normal machines is reversed. This limits the optimization space if all normal machines are full, which corresponds to the analysis in Section 5.3.
As shown in Table 7, when compared with heterogeneous algorithms, WindGP also shows >1.5×>1.5$\times$> 1.5 × and >1.2×>1.2$\times$> 1.2 × speedup in the running time of PageRank and triangle counting respectively, as it performs collaborative optimization on machines with heterogeneous memory, computing power and network traffic. In contrast, (Verma and Zeng, 2005) only optimize load balance while (Mayer et al., 2016) targets at communication cost. Thus, the computing time of (Verma and Zeng, 2005) is similar to ours, but its communication time is $\sim$50longer. (Mayer et al., 2016) has ¿20% longer computing time, and its communication time is also $\sim$18longer as it lacks the best-first search and comprehensive subgraph-local tuning. Besides, none of them take care of the heterogeneous memory capacity and they have ¿5% longer executing time. More experimental details are omitted due to space limit, which can be supplemented in the full version if required.
@@
Table 6. Performance of distributed graph computing (unit:s)
Data PageRank SSSP
HDRF NE WindGP HDRF NE WindGP
TW 2,380 4,417 353 1,046 1,820 182
DB xx xx xx xx xx xx
FR xx xx xx xx xx xx
YH xx xx xx xx xx xx
@@
Data PageRank SSSP
@@
DB xx xx xx xx xx xx
FR xx xx xx xx xx xx
YH xx xx xx xx xx xx

### 06 Mar 2024 v2

@@ 5.2. Evaluation by the metric $TC$
Analysis of optimization techniques. Recall that WindGP propose three novel techniques: preprocessing capacity, best-first search, post-processing. To evaluate the efficiency of each technique in WindGP, we conduct the experiments on six real graphs and depict the results in Figure 8. Let WindGP be the entire solution, ${}^{+}$ denotes the solution that removes post-processing from WindGP. Similarly, $*{}^{*}$ denotes the solution that removes best-first search from ${}^{+}$, while ${}^{-}$ is the naive solution without all optimization techniques, which iteratively explores the graph to form a partition as long as the machine memory and load balance ratio permit. Its expansion scheme is similar to NE (Zhang et al., 2017), which tends to form a subgraph with high cohesion by connectivity-based expansion. In order to view the details of all partitions precisely, the histograms of the costs of different partitions on CP and LJ are also listed in Figure 9, Figure 10 and Figure 11, respectively.
@@
Obviously, $TC$ of $*{}^{*}$ is much lower than ${}^{-}$, and the speedup is 5×5$\times$5 × on TW, 4.5×4.5$\times$4.5 × on CO, $\times3$ on LJ, 3.8×3.8$\times$3.8 × on PO, 7.5×7.5$\times$7.5 × on CP, and $\times8$ on RN, respectively. This prominent improvement comes from the pre-computed capacity of edges in each machine. Different from the naive solution that only considers the upper bound of machine memory and homogeneous load balance ratio, our strategy combines the computation cost and the memory size of all machines to generate a nearly best plan that can balance the computation cost across these heterogeneous machines (as shown in Figure 9(b) and Figure 10(b)). Note that traditional partition methods use $\frac{|E|}{p}$ as the capacity of edges, which can not be used in heterogeneous scenarios because the memory of some machines may be not enough while other machines may have memory size larger than $\frac{|E|}{p}$. Besides, the traditional threshold does not distinguish the difference of node/edge computation cost between heterogeneous machines. Recall that in Section 3.2 we deduce the error bound $\frac{p^{2}}{|E|}$, and the experiment on small graphs with hundreds of edges shows that the difference between our solution and the optimal is always within 5%, which can be further refined by the post-processing. According to the theoretical bound, on larger graphs the error should be much smaller.
@@
The second technique (i.e., the best-first search) further boosts the performance by optimizing the expansion process, leading to >1.1×>1.1$\times$> 1.1 × on TW, CO and LJ. The performance gain is mainly acquired by lowering the total communication cost (as shown in Figure 9(c) and Figure 10(c)). The degree balanced generation as well as border generation can help reduce the number of cross-machine vertices and improve the subgraph cohesion. However, on low-degree graphs (CP and RN), the speedup is rather limited (only 1.04$\sim$ 1.04$\times$ 1.04 ×) due to the small percentage of communication cost. On CP, the percentage of communication cost is $<30\%$ on sparse graphs (i.e., graphs with low average degree $\frac{|E|}{|V|}$), while it is $>50\%$ on LJ and CO. Especially, the effect is tiny on RN because it is mesh-like graph and its structure is naturally balanced.
@@
Finally, the post-processing technique (subgraph-local search) brings >1.15×>1.15$\times$> 1.15 × speedup by flattening the total cost between all machines. Note that the previous two techniques balance the computation cost and the communication cost respectively, however, they do not balance the total cost well. In fact, the effect of best-first search is sometimes restricted because it can not disturb the balance of computation cost. Besides, during expansion the information is not complete (e.g., the first partition has no communication cost), which further drags down the performance. Subgraph-local search breaks this restriction by allowing the computation cost to be imbalanced between heterogeneous machines, as long as the total cost drops. Based on the global view with complete information, it moves or swaps edges between machines to lower the highest total cost.
@@
Comparison with counterparts. Figure 12 lists the comparison of WindGP (the entire solution) with the state-of-the-art algorithms on heterogeneous machines. As reported in previous work (Zhang et al., 2017), METIS can not partition too large graphs like TW due to its prohibitive memory occupation. Among four counterparts, NE performs the best on most datasets except for CP and RN. The two low-degree graphs have relatively larger computation cost and lower communication cost, thus the expansion scheme of NE can not optimize a lot. In contrast, METIS and HDRF perform better on CP and RN due to their elaborated consideration of computation cost in the partition objective. As for EBV solution, though it is carefully designed to reduce the impact of skew on power-law graphs, its objective is not equivalent to the maximum total cost and it can not be well adapted to heterogeneous cases.
@@
Due to the lightweight preprocessing and limited iterations in post-precessing, the complexity of WindGP is linear to the graph size and machine number, which is similar to NE. As for the executing time of partition methods, Table 11 shows that their executing time is in the same scale, while WindGP is 11% slower than NE, which is acceptable in real cases as partitioning is not time sensitive. Besides, extra experiments (Table 10) with the best two counterparts (HDRF and NE) show that the performance of WindGP on homogeneous clusters is not worser than others, as Section 2.1 proves that the $TC$ metric is equivalent to load balance and $RF$ in homogeneous scenarios. Furthermore, the running time of distributed graph algorithms on the partition results is also evaluated (see Section 5.4), which corresponds to the comparative results of $TC$.
Table 10. Evaluation of PageRank on homogeneous 30-machine cluster on LJ
Alg. $\alpha^{\prime}$ $RF$ $TC$ time (s)
HDRF 1.1 3.33 105M 92
NE 1.1 1.55 19M 18
WindGP 1.0 1.56 20M 18
Table 11. Evaluation of partitioning time (s) on traditional methods
Dataset METIS HDRF NE EBV WindGP
CO 200 15 80 91 89
LJ 71 5 23 30 25
PO 68 4 21 24 23
CP 32 3 12 14 13
RN 6 1 2 2.3 2.1
@@ 5.3. Scalability Test
Scalability with the number of machine types. In previous experiments, we use two kinds of machines by default. Here we vary the number of machine types from one to six and test the performance on LJ with 30 machines. The added type is extracted from normal machines, increase the machine memory, computation and communication cost. Note that the 2-type setting here is different from the default setting. Compared to the single-type setting (i.e., homogeneous cases where all machines are normal machine), 2-type setting transforms 5 machines to slightly bigger machine with larger computation and communication cost as well as larger memory size.
Generally, $TC$ increases as the number of machine types grows, i.e., the homogeneous cases achieve the minimum $TC$ for all solutions. Note that NE performs extremely good on homogeneous cases, even better than WindGP. However, both NE and EBV fail to adapt to heterogeneous cases and their $TC$ rises sharply when the number grows. Comparatively, WindGP achieves the slowest growth due to its flexible preprocessing technique, which can pre-computes appropriate edge capacities for all machine settings. Thus, WindGP is not sensitive to the number of machine types.
@@
DataSet PageRank SSSP
(Verma and Zeng, 2005) GrapH HaSGP HAEP WindGP speedup (Verma and Zeng, 2005) GrapH HaSGP HAEP WindGP speedup
TW 1287 681 1,015 529 353 1.49$\times$ 667 309 618 254 182 1.39$\times$
DB 1168 609 970 501 280 1.78$\times$ 621 305 597 226 158 1.43$\times$
FR 2,017 2,259 2,380 1,307 681 1.92$\times$ 1,132 1,218 1,846 580 312 1.86$\times$
YH 3,204 3,980 3,012 2,201 1,048 2.1$\times$ 1.910 2,203 2,018 1,038 509 2.04$\times$
@@ 5.4. Evaluation of Distributed Graph Algorithms
To evaluate the real performance of distributed graph computing, we choose PageRank (Pretto, 2002) and Single-Source Shortest Path ((Forster and Nanongkai, 2017)) and perform the test on Plato system (Tencent, 2019). These two algorithms are the most representative dense and sparse algorithms, respectively. In PageRank, all edges are computed and all nodes are updated in each iteration. In contrast, the number of active nodes as well as edges grows first until the largest point, then reduces until the end. Modern machines have large memory and computing power, thus graphs with billions of edges are added in this section. Except for TW, referring to (Zhong et al., 2023) and (Zeng et al., 2022c), we add DB (0.23B nodes and 1.1B edges), FR (65M nodes, 1.8B edges), and YH (0.41B nodes and 2.8B edges). Similar to (Zhang et al., 2017), we use nine machines connected with Gigabit Ethernet, and $\frac{1}{3}$ of them are super machines. The resource of the real cluster is listed below:
  • 3 super machines: 1.6GHz (4), 6GB memory, 100Gbps
  • 6 normal machines: 1.6GHz (8), 2GB memory, 150Gbps
The memory capacity, computation cost and communication cost of each machine ($M_{i}$, $C_{i}^{node}$, $C_{i}^{edge}$ and $C_{i}^{com}$) is estimated according to the quantification process in Section 2.1. Note that the machine configuration is different from previous settings, thus $TC$ is also evaluated. Table 14, 15, 16 and 17 list the comparison of $TC$ and distributed running time (seconds).
@@
CO 938,849,480 565,506,575 189,439,055
LJ 182,987,265 205,252,940 57,918,640
PO 278,052,160 196,995,825 50,954,100
CP 106,732,065 195,154,235 31,504,615
RN 9,632,230 61,151,165 4,950,860
Table 15. Performance of distributed graph computing (unit:s)
Data PageRank TriangleCount
HDRF NE WindGP HDRF NE WindGP
TW 2,380 4,417 353 1,046 1,520 182
CO 791 398 125 325 147 97
LJ 112 201 54 33 33 25
PO 130 187 41 29 22 18
CP 64 166 27 11 12 11
RN 11 52 9 7 9 5
Table 16. Performance of distributed graph computing (unit:s)
DataSet $TC$ PageRank SSSP
HDRF NE WindGP HDRF NE WindGP HDRF NE WindGP
TW 2.7G 5.6G 0.4G 2,380 4,417 353 1,046 1,820 182
DB 2.5G 4.2G 0.3G 2,146 3,709 280 965 1,227 158
FR 5.1G 4.4G 1.0G 5,012 4,183 681 1,398 1,260 312
YH 6.8G 5.9G 1.7G 8,149 5,980 1,048 2,012 1,865 509
@@
Comparison with Heterogeneous Solutions. As shown in Table 13, when compared with heterogeneous solutions ((Verma and Zeng, 2005), GrapH, HaSGP, HAEP), WindGP also shows >1.49×>1.49$\times$> 1.49 × and >1.39×>1.39$\times$> 1.39 × speedup in the running time of PageRank and SSSP respectively, as it performs collaborative optimization on machines with heterogeneous memory capacity, computing power and network bandwidth. Note that the speedup is calculated by comparing WindGP with the best counterpart HAEP. None of the counterparts takes care of the heterogeneous memory capacity, thus they can not compete with WindGP when optimizing the computation cost. Specifically, (Verma and Zeng, 2005) only optimizes load balance while (Mayer et al., 2016) targets at communication cost. Thus, the computing time of (Verma and Zeng, 2005) is similar to ours, but its communication time is $\sim$50longer. (Mayer et al., 2016) has ¿20% longer computing time, and its communication time is also $\sim$18longer as it lacks the best-first search and comprehensive subgraph-local tuning. HaSGP is a streaming partition algorithm, which performs terrible when running on severely skewed graphs like TW and DB. The degree distribution of FR and YH is much more balanced, thus (Verma and Zeng, 2005) performs better than GrapH as it balances the computation cost better while existing communication-optimized methods (NE, GrapH, HAEP) have non-prominent effect on these two graphs. Inherently, most nodes have the same scale of neighbors, and there is no significant difference when we select which one as the boundary. However, WindGP can optimize this case well because the border generation in Section 3.3 restricts the border vertex to reside in as fewer machines as possible.
@@
Data PageRank TriangleCount
@@
CO 411 208 125 198 116 97
LJ 98 81 54 32 31 25
PO 201 126 41 38 27 18
CP 53 108 27 18 20 11
RN 10 32 9 7 21 5
Comparing PageRank and SSSP, obviously the speedup on PageRank is always higher than that on SSSP. The biggest difference is on DB, which has the largest degree skew and the lowest averaged degree (i.e., 4.7) among four large graphs. In PageRank, all nodes need to join each iteration while only partial nodes are active in SSSP. The largest skew enlarges the optimization space of WindGP in PageRank, while the lowest averaged degree limits our techniques in SSSP (just like the cases in mesh-like RN). This indicates that WindGP works well on dense algorithms like PageRank, but further improvement can be developed on sparse algorithms like SSSP.
[…]
Table 18. Evaluation of partitioning time on heterogeneous methods
Dataset (Verma and Zeng, 2005) GrapH HaSGP HAEP WindGP
TW 105 92 91 112 101
DB 109 90 90 110 108
FR 156 137 138 170 152
YH 301 218 210 312 293